Draw Free Body Diagrams (FBD) for member be and ce 12 kN/m 6 kN 4 m V. -30 kN H6 kN M 78 kN -m v,-42 kN
Answer #1

Solution: าว่า 12 KN rn d. G KN Ma 78 KN. V 42 KN 4.SKN m 37.S kN.m 3okN 4SKN 6 kN 73 KN GK N 42 KNBC 12. 33 kM 20 3 3 4-2 KN AC Just below e ( Mc=Jo KN.m) [(It) 6KN 6 KN 4-2 kN=6- 6 O kN de (2-2) KN/m 6 KN/m d. 3oKN 2. 21de 12 kN/m 6 KM -ใ 30 KN 12t3